# EMGauss: Continuous Slice-to-3D Reconstruction via Dynamic Gaussian Modeling in Volume Electron Microscopy

### v1

@@ 1 Introduction
Volume electron microscopy (vEM) has revolutionized the understanding of biological ultrastructures, enabling nanoscale 3D imaging of cells, tissues, and even entire organisms [collinson2023volume, glausier2025volume, li2025situ, robert2025improving, turegano2024tracing]. However, due to the inherent trade-off among resolution, field of view, and acquisition time—commonly referred to as the “impossible triangle”—direct acquisition of isotropic data remains costly and inefficient [heinrich2017deep, peddie2022volume]. This often results in anisotropic volumes, with axial (z) resolution substantially lower than the in-plane (xy) resolution.
Figure 1: Different paradigms of isotropic reconstruction from anisotropic datasets. Existing approaches typically extract volume features from abundant xy-plane images. After stacking anisotropic slices along the z-axis (top-left), they: (a) Video interpolation along the z-axis: models are trained along the x or y direction, where adjacent frames are input to synthesize the intermediate frame; during inference, the trained model is repurposed to interpolate along the z-axis, generating intermediate slices corresponding to continuous temporal coordinates ($\tau\in[0,1]$). (b) Image super-resolution on xz/yz planes: Since the xy-plane images are of much higher resolution, a low-resolution counterpart can be generated either by manual downsampling [emdiffuse] or by learning a degradation model [cycleganir], forming paired HR–LR supervision for training; during inference, the resulting super-resolution model is then applied along the z-axis to enhance xz/yz slices. Both paradigms implicitly assume spatial isotropy. In contrast, our method directly performs inference in continuous 3D space, without relying on such isotropic assumptions.
@@
Deep learning has emerged as a powerful tool to address this anisotropy by computationally enhancing axial resolution [heinrich2017deep, emdiffuse, cycleganir, hagita2018super, he2023isovem]. These methods aim to restore high-resolution details from inexpensive anisotropic acquisitions, thereby boosting imaging throughput while preserving structural fidelity. Conventional learning-based approaches typically rely on supervised training using paired high- and low-resolution volumes [heinrich2017deep], or on partially isotropic samples [emdiffuse], leading to strong dependence on well-curated data and poor adaptability across different domains and imaging conditions. In practice, isotropic reference data are rarely available, as the need for isotropic reconstruction mostly arises from the absence of such acquisitions. Self-supervised alternatives have been explored to alleviate the data reliance on isotropic data, including GAN-based [cycleganir, hagita2018super] and diffusion-based [pan2023diffuseir, emdiffuse] methods. Their key insight is to leverage the high-resolution xy-planes as an internal reference. As shown in Figure 1, to reconstruct a volume from anisotropic slices, they either perform video frame-interpolation between consecutive xy-slices to generate intermediate frames [cycleganir, he2023isovem], or apply image super-resolution to orthogonal views (e.g., xz or yz) to improve resolution along the z-axis [emdiffuse]. Both strategies rely on the critical assumption that local tissue structures are approximately isotropic across the x, y, and z dimensions, such that geometric patterns in the high-resolution xy-planes can be extrapolated along the axial direction.
@@
In practice, however, the assumption of isotropy is often violated in real biological specimens, where morphological anisotropy is prevalent—e.g., elongated neuronal fibers or dendritic spines exhibit strong directional continuity [zhang2012noddi, beaulieu2002basis]. Consequently, these methods suffer from limitations in accurately segmenting or analyzing such structures. This motivates the need for frameworks that move beyond 2D image priors to directly reason about 3D space and structural continuity, capable of preserving the true morphology of biological specimens.
@@
In this way, our method fundamentally circumvents the domain mismatch issue inherent in prior isotropic-assumed models. Moreover, EMGauss offers two additional advantages. First, unlike methods such as EMDiffuse [emdiffuse], which require additional isotropic sub-volumes of the same tissue for training, our approach operates in a fully self-contained optimization loop that uses only the anisotropic slices from the target volume. This makes it suitable for data-scarce or computationally constrained scenarios without the need for large-scale pretraining or auxiliary datasets. Second, by explicitly modeling geometric and photometric evolution across slices, our method enforces spatial continuity and physical coherence in the reconstructed volume, leading to stable training and reduced artifacts in the final reconstructed volume.
@@ 2.1 Isotropic Reconstruction of Volume EM
Volume Electron Microscopy (vEM) refers to a suite of techniques for generating 3D volumes of biological tissues, which are foundational for analyzing complex 3D cellular structures [li2025situ, glausier2025volume, turegano2024tracing, zhao2024application]. A standard vEM workflow consists of sequentially sectioning and imaging a specimen, then computationally aligning and reconstructing them into a coherent 3D volume [peddie2022volume, collinson2023volume]. However, achieving isotropic reconstruction for large volumes remains elusive, as high-throughput methods like ssTEM [briggman2006towards],ssSEM [hayworth2014imaging] and SBF-SEM [denk2004serial] prioritize speed over Z-axis fidelity, while isotropic techniques like enhanced FIB-SEM [bushby2011imaging] are limited to small volumes and require specialized instrumentation [peddie2022volume]. Thus, achieving isotropic reconstruction for biological specimens remains a persistent challenge.
@@
Early approaches focused on slice interpolation or optical-flow-based alignment [carata2011improving, gonzalez2022optical, ferede2025z, joshi2025interpolai]. While straightforward to implement, they often introduce substantial artifacts and errors under high anisotropy factors. Deep learning-based approaches are shifting from fully supervised learning to unsupervised learning. Fully supervised methods such as 3DSRUNet [heinrich2017deep] employed paired high- and low-resolution volumes for training, demonstrating strong interpolation quality on synthetic data. However, collecting such paired isotropicanisotropic samples from the same tissue imposes stringent hardware and imaging requirements—often necessitating dual-mode acquisition or repeated scans under different sectioning protocols—which are impractical for most vEM systems [peddie2022volume]. On the other hand, synthetic degradations (e.g., row removal or Gaussian blur) poorly approximate real imaging processes, leading to domain gaps and limited generalization on real vEM data. Consequently, self-supervised methods that do not require isotropic training datasets are more suitable for isotropic reconstruction in vEM.
@@
Pioneering self-supervised works [weigert2018content, krull2019noise2void] generate pseudo HR-LR pairs internally, for instance, by masking or reinterpreting image axes, yet these methods typically assume a known point spread function (PSF). In practice, the degradation process is unknown and often deviates significantly from handcrafted models, limiting their generalization in real applications. Alternative methods like CycleGAN [cycleganir] rely only on target tissue data to learn a degradation model, but its adversarial formulation often leads to instability and artifacts such as texture aliasing or geometric distortion. More recently, diffusion-based models [emdiffuse, pan2023diffuseir] learn generative priors from abundant xy-plane images, using them to enhance axial resolution. While they achieve higher fidelity and stability than GAN-based methods, these 2D slice-wise approaches still assume isotropy across spatial dimensions and suffer from domain mismatch when transferring degradations learned from lateral (xy-planes) to the axial (z-axis) direction. Our work addresses this limitation by directly modeling continuous 3D geometry across slices, rather than relying on isotropy assumptions or proxy 2D priors.
@@
Our work builds upon the recent advancements in real-time neural rendering, particularly the paradigm introduced by 3D Gaussian Splatting (3DGS) and its subsequent extensions to dynamic scenes. 3DGS [kerbl20233d] represents scenes as a collection of anisotropic 3D Gaussians, each defined by position, covariance, opacity, and spherical harmonic coefficients for view-dependent color. This formulation enables rapid optimization and fast rendering at over 100 FPS at 1080p resolution, significantly outperforming prior explicit [muller2022instant] and implicit [mildenhall2021nerf] representations in both speed and quality.
To extend 3DGS to dynamic scenes, several variants introduce temporal modeling through deformation fields that warp canonical Gaussians to observed timestamps. For instance, Deformable 3D Gaussians [yang2024deformable] employs a lightweight multi-layer perceptron (MLP) to predict per-Gaussian deformations from input timestamps, enabling high-fidelity monocular reconstruction of non-rigid motions. In contrast, 4D Gaussian Splatting (4DGS) [wu20244d] constructs a deformation field using a 4D neural voxel representation decomposed via a HexPlane-inspired encoding, followed by a compact MLP decoder for efficient motion prediction. These methods collectively advance real-time dynamic scene reconstruction by leveraging the efficiency of Gaussian primitives while addressing motion modeling challenges. In parallel, there is an emerging line of work that applies 2D Gaussian Splatting) for image restoration, compression, and other 2D image–space tasks [zhang20242d, zhang2024gaussianimage, hu2025gaussiansr, zhu2025large]. These 2D‐space extensions reveal the broader applicability of Gaussian splatting primitives in imaging tasks beyond volumetric reconstruction.
@@ 3.1 Preliminaries
Our method builds upon Deformable 3D Gaussians [yang2024deformable],an extension of 3D Gaussian Splatting (3DGS) [kerbl20233d] for dynamic representations. 3DGS models a scene as a collection of explicit 3D Gaussian primitives, each parameterized by an opacity $o\in[0,1]$, a center $\mu\in\mathbb{R}^{3\times 1}$, and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$:
@@
Building on this foundation, Deformable 3D Gaussians [yang2024deformable] extend 3DGS to dynamic scenes by learning a canonical Gaussian set $\mathcal{G}c={(\mu_{i},S_{i},q_{i},o_{i},\mathcal{C}i)}_{i=1}^{M}$ along with a lightweight MLP-based deformation network $\Phi\theta$. Conditioned on a temporal coordinate $t\in[0,1]$, $\Phi_{\theta}$ predicts per-Gaussian offsets:
@@
Figure 3: Isotropic xy-slice reconstruction results on EPFL dataset [lucchi2013learning].
@@ 3.2 Slice-to-3D Dynamic Gaussians
As shown in Figure 2, from a canonical Gaussian set $\mathcal{G}c={(\mu_{i},S_{i},R_{i},o_{i},\mathcal{C}i)}_{i=1}^{M}$ initially optimized from the observed anisotropic frames, the deformation network $\Phi_{\theta}$ predicts local offsets for each Gaussian as:
@@ 3.3 Bootstrapping with Pseudo Gaussian Targets
Figure 4: Isotropic xy-slice reconstruction results on FIB-25 dataset [takemura2015synaptic].
@@ 3.4 Training Pipeline
Our model can effectively function with only RGB photometric supervision. Following 3DGS [kerbl20233d], we compute the deviation between the rendered image and the ground-truth RGB frame using an $\ell_{1}$ loss combined with a D-SSIM regularization term. For pseudo-supervised iterations, we apply the same objective but scale the loss weight $\lambda_{\text{pseudo}}$ by an exponentially increasing factor, gradually enhancing the impact of pseudo supervision as training progresses.
@@
We implement EMGauss in PyTorch [paszke2019pytorch] and optimize with Adam [kingma2014adam] following the default hyperparameters of 3DGS [kerbl20233d] for Gaussian attributes and learning rates. Training proceeds in three phases: a warm-up phase of 22k iterations optimizing the canonical Gaussian set $\mathcal{G}_{c}$ with the deformation MLP frozen, a short joint-training phase of 11k iterations enabling the $\mathcal{G}_{c}$ and the deformation MLP, and a final phase of $15k$ iterations with the EMA teacher providing pseudo supervision on unseen timestamps until convergence. During the pseudo-supervised phase, the corresponding loss weight is linearly ramped from $0.1$ to $1.0$ between 33k and 1010k iterations. The EMA teacher employs a decay factor of $\alpha=0.995$. All experiments are conducted on an NVIDIA RTX 3090 GPU.
@@ 4.1 Experimental Settings
We conduct experiments on a diverse set of large-scale electron microscopy (EM) datasets covering both isotropic and anisotropic imaging conditions, as shown in Table 1. For isotropic data, we select the EPFL mouse brain dataset [lucchi2013learning] acquired from the CA1 hippocampus region of the mouse brain with a uniform voxel resolution of $5\times 5\times 5$ nm, and the FIB-25 drosophila brain dataset [takemura2015synaptic] with a resolution of $8\times 8\times 8$ nm. Both datasets are collected by FIB-SEM. To simulate anisotropic imaging conditions and enable controlled evaluation, we subsample the z-dimension of these isotropic volumes by factors of $\times 6$, treating the withheld intermediate slices as ground truth for quantitative assessment of reconstruction quality. For each dataset, we follow CycleGAN-IR [cycleganir] to use a sub-volume of $500\times 500\times 500$ pixels for our experiments. To further assess performance under real-world anisotropic acquisition, we evaluate on the FANC dataset [phelps2021reconstruction], reconstructed from the ventral nerve cord of an adult female Drosophila with an anisotropy ratio of $10$ ($4\times 4\times 40$ nm). For the anisotropic dataset, we use a sub-volume of $500\times 500\times 50$ for our experiments. These datasets collectively span multiple organisms, tissue types, and imaging resolutions, providing a robust benchmark for both simulated and real anisotropic reconstruction. We use PSNR, SSIM [wang2004image], and FSIM [zhang2011fsim] to evaluate all the methods.
Datasets Isotropy Resolution (nm) Anisotropy ratio Issue
EPFL [lucchi2013learning] $\checkmark$ $5\times 5\times 5$ $6\times$ brain
FIB-25 [takemura2015synaptic] $\checkmark$ $8\times 8\times 8$ $6\times$ brain
FANC [phelps2021reconstruction] $\times$ $4\times 4\times 40$ $10\times$ nerve
Table 2: Training cost for generating one sample. To effectively generate the in-between slices, we only need a small sub-batch of the overall data, achieving the minimum requirement on data amount and time consumption.
Methods Time Frames
CycleGAN-IR[cycleganir] 2h 500
EMDiffuse[emdiffuse] 24h 128
Ours 0.5h 6
@@
We compare our method against several state-of-the-art methods for isotropic reconstruction from anisotropic EM data. The baseline methods include: (1) vEMDiffuse-a from EMDiffuse [emdiffuse], a diffusion-based model that requires additional training data from the same tissue type beyond the target dataset to learn anisotropic-to-isotropic mapping. To train the model, we use another volume from the same tissue as the target volume, with the same spatial size, as the training data. (2) CycleGAN-IR [cycleganir], an unsupervised framework using a CycleGAN [zhu2017unpaired] architecture that learns cross-domain mappings between unpaired high-resolution (HR) lateral and low-resolution (LR) axial slices through cycle-consistent adversarial training. For vEMDiffuse-a and CycleGAN-IR, we train with the default hyperparameters in their codebases. In Table 2 we compare the training cost for generating one sample by each method.

### v2

@@ 1 Introduction
Volume electron microscopy (vEM) has revolutionized the understanding of biological ultrastructures, enabling nanoscale 3D imaging of cells, tissues, and even entire organisms [5, 9, 20, 30, 32]. However, due to the inherent trade-off among resolution, field of view, and acquisition time—commonly referred to as the “impossible triangle”—direct acquisition of isotropic data remains costly and inefficient [14, 27]. This often results in anisotropic volumes, with axial (z) resolution substantially lower than the in-plane (xy) resolution.
@@
Deep learning has emerged as a powerful tool to address this anisotropy by computationally enhancing axial resolution [14, 21, 6, 11, 13]. These methods aim to restore high-resolution details from inexpensive anisotropic acquisitions, thereby boosting imaging throughput while preserving structural fidelity. Conventional learning-based approaches typically rely on supervised training using paired high- and low-resolution volumes [14], or on partially isotropic samples [21], leading to strong dependence on well-curated data and poor adaptability across different domains and imaging conditions. In practice, isotropic reference data are rarely available, as the need for isotropic reconstruction mostly arises from the absence of such acquisitions. Self-supervised alternatives have been explored to alleviate the data reliance on isotropic data, including GAN-based [6, 11] and diffusion-based [25, 21] methods. Their key insight is to leverage the high-resolution xy-planes as an internal reference. As shown in Figure 1, to reconstruct a volume from anisotropic slices, they either perform video frame-interpolation between consecutive xy-slices to generate intermediate frames [6, 13], or apply image super-resolution to orthogonal views (e.g., xz or yz) to improve resolution along the z-axis [21]. Both strategies rely on the critical assumption that local tissue structures are approximately isotropic across the x, y, and z dimensions, such that geometric patterns in the high-resolution xy-planes can be extrapolated along the axial direction.
@@
In practice, however, the assumption of isotropy is often violated in real biological specimens, where morphological anisotropy is prevalent—e.g., elongated neuronal fibers or dendritic spines exhibit strong directional continuity [37, 1]. Consequently, these methods suffer from limitations in accurately segmenting or analyzing such structures. This motivates the need for frameworks that move beyond 2D image priors to directly reason about 3D space and structural continuity, capable of preserving the true morphology of biological specimens.
Figure 1: Different paradigms of isotropic reconstruction from anisotropic datasets. Existing approaches typically extract volume features from abundant xy-plane images. After stacking anisotropic slices along the z-axis (top), they: (a) Video interpolation along the z-axis: models are trained along the x or y direction, where adjacent frames are input to synthesize the intermediate frame; during inference, the trained model is repurposed to interpolate along the z-axis, generating intermediate slices corresponding to continuous temporal coordinates ($\tau\in[0,1]$). (b) Image super-resolution on xz/yz planes: Since the xy-plane images are of much higher resolution, a low-resolution counterpart can be generated either by manual downsampling [21] or by learning a degradation model [6], forming paired HR–LR supervision for training; during inference, the resulting super-resolution model is then applied along the z-axis to enhance xz/yz slices. Both paradigms implicitly assume spatial isotropy. In contrast, our method directly performs inference in continuous 3D space, without relying on such isotropic assumptions.
@@
In this way, our method fundamentally circumvents the domain mismatch issue inherent in prior isotropic-assumed models. Moreover, EMGauss offers two additional advantages. First, unlike methods such as EMDiffuse [21], which require additional isotropic sub-volumes of the same tissue for training, our approach operates in a fully self-contained optimization loop that uses only the anisotropic slices from the target volume. This makes it suitable for data-scarce or computationally constrained scenarios without the need for large-scale pretraining or auxiliary datasets. Second, by explicitly modeling geometric and photometric evolution across slices, our method enforces spatial continuity and physical coherence in the reconstructed volume, leading to stable training and reduced artifacts in the final reconstructed volume.
@@ 2.1 Isotropic Reconstruction of Volume EM
Volume Electron Microscopy (vEM) refers to a suite of techniques for generating 3D volumes of biological tissues, which are foundational for analyzing complex 3D cellular structures [20, 9, 32, 41]. A standard vEM workflow consists of sequentially sectioning and imaging a specimen, then computationally aligning and reconstructing them into a coherent 3D volume [27, 5]. However, achieving isotropic reconstruction for large volumes remains elusive, as high-throughput methods like ssTEM [2],ssSEM [12] and SBF-SEM [7] prioritize speed over Z-axis fidelity, while isotropic techniques like enhanced FIB-SEM [3] are limited to small volumes and require specialized instrumentation [27]. Thus, achieving isotropic reconstruction for biological specimens remains a persistent challenge.
@@
Early approaches addressed anisotropic reconstruction through slice interpolation or optical-flow-based alignment [4, 10, 8, 16]. While straightforward to implement, these methods often introduce severe artifacts under high anisotropy factors. Deep learning approaches initially relied on fully supervised training. For instance, 3DSRUNet [14] learns interpolation from paired high- and low-resolution volumes and achieves strong performance on synthetic data. However, acquiring such paired isotropic/anisotropic volumes from the same tissue requires specialized imaging protocols, such as dual-mode acquisition or repeated scans, which are rarely feasible for most vEM systems [27]. Moreover, synthetic degradations (e.g., row removal or Gaussian blur) poorly approximate real imaging processes, leading to domain gaps and limited generalization. These limitations motivate self-supervised approaches that avoid the need for isotropic training data.
@@
Early self-supervised methods [34, 19] construct pseudo HR–LR pairs internally, for example by masking pixels or exploiting axis re-interpretation. However, they typically assume a known point spread function (PSF), whereas the true degradation in practice is often unknown and deviates from handcrafted models. Alternative strategies such as CycleGAN [6] attempt to learn the degradation process directly from target tissue data. However, adversarial training can introduce instability and artifacts, including texture aliasing and geometric distortion. More recently, diffusion-based models [21, 25] learn generative priors from abundant xy-plane images and leverage them to enhance axial resolution. Although these approaches improve fidelity and stability compared with GAN-based methods, they operate in a slice-wise 2D manner and implicitly assume spatial isotropy. As a result, degradations learned in the lateral (xy) plane may not transfer well to the axial (z) direction. In contrast, our method directly models continuous 3D geometry across slices, avoiding isotropy assumptions and enabling more faithful reconstruction of anisotropic structures.
@@
Recent advances in neural rendering have been significantly driven by the emergence of 3D Gaussian Splatting (3DGS) and its extensions to dynamic scenes. 3DGS [17] represents a scene as a set of anisotropic 3D Gaussians parameterized by position, covariance, opacity, and spherical harmonic coefficients for view-dependent appearance. This explicit representation enables efficient optimization and high-quality real-time rendering, substantially improving both speed and fidelity compared with earlier explicit [24] and implicit [23] scene representations.
To handle dynamic scenes, subsequent works introduce temporal modeling through deformation fields that map canonical Gaussians to time-dependent observations. Deformable 3D Gaussians [36], for example, employ a lightweight multi-layer perceptron (MLP) to predict per-Gaussian deformations conditioned on timestamps, enabling reconstruction of non-rigid motions from monocular videos. Similarly, 4D Gaussian Splatting (4DGS) [35] models motion using a 4D neural voxel representation decomposed with a HexPlane-style encoding, followed by a compact MLP decoder for efficient deformation prediction. These approaches extend the efficiency of Gaussian primitives to dynamic scene reconstruction while explicitly modeling temporal variation.
Beyond volumetric rendering, recent studies also explore Gaussian splatting in the 2D image domain for tasks such as image restoration, compression, and super-resolution [39, 40, 15, 43]. These works highlight the flexibility of Gaussian primitives and suggest their broader applicability to imaging problems beyond 3D scene reconstruction.
@@ 3.1 Preliminaries
Our method builds upon Deformable 3D Gaussians [36],an extension of 3D Gaussian Splatting (3DGS) [17] for dynamic representations. 3DGS models a scene as a collection of explicit 3D Gaussian primitives, each parameterized by an opacity $o\in[0,1]$, a center $\mu\in\mathbb{R}^{3\times 1}$, and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$:
@@
Building on this foundation, Deformable 3D Gaussians [36] extend 3DGS to dynamic scenes by learning a canonical Gaussian set $\mathcal{G}c={(\mu_{i},S_{i},q_{i},o_{i},\mathcal{C}i)}_{i=1}^{M}$ along with a lightweight MLP-based deformation network $\Phi\theta$. Conditioned on a temporal coordinate $t\in[0,1]$, $\Phi_{\theta}$ predicts per-Gaussian offsets:
@@
Figure 3: Isotropic xy-slice reconstruction results on EPFL dataset [22].
@@ 3.2 Slice-to-3D Dynamic Gaussians
As shown in Figure 2, starting from a canonical Gaussian set $\mathcal{G}c={(\mu_{i},S_{i},R_{i},o_{i},\mathcal{C}i)}_{i=1}^{M}$ initially optimized from the observed anisotropic frames, the deformation network $\Phi_{\theta}$ predicts local offsets for each Gaussian as:
@@ 3.3 Bootstrapping with Pseudo Gaussian Targets
Figure 4: Isotropic xy-slice reconstruction results on FIB-25 dataset [31].
@@ 3.4 Training Pipeline
Our model can effectively function with only RGB photometric supervision. Following 3DGS [17], we compute the deviation between the rendered image and the ground-truth RGB frame using an $\ell_{1}$ loss combined with a D-SSIM regularization term. For pseudo-supervised iterations, we apply the same objective but scale the loss weight $\lambda_{\text{pseudo}}$ by an exponentially increasing factor, gradually enhancing the impact of pseudo supervision as training progresses.
@@
We implement EMGauss in PyTorch [26] and optimize with Adam [18] following the default hyperparameters of 3DGS [17] for Gaussian attributes and learning rates. Training proceeds in three phases: a warm-up phase of 22k iterations optimizing the canonical Gaussian set $\mathcal{G}_{c}$ with the deformation MLP frozen, a short joint-training phase of 11k iterations enabling the $\mathcal{G}_{c}$ and the deformation MLP, and a final phase of $15k$ iterations with the EMA teacher providing pseudo supervision on unseen timestamps until convergence. During the pseudo-supervised phase, the corresponding loss weight is linearly ramped from $0.1$ to $1.0$ between 33k and 1010k iterations. The EMA teacher employs a decay factor of $\alpha=0.995$. All experiments are conducted on an NVIDIA RTX 3090 GPU.
@@ 4.1 Experimental Settings
We conduct experiments on diverse large-scale electron microscopy (EM) datasets under both isotropic and anisotropic imaging conditions, as shown in Table 1. For isotropic data, we select the EPFL mouse brain dataset [22] acquired from the CA1 hippocampus region of the mouse brain with a uniform voxel resolution of $5\times 5\times 5$ nm, and the FIB-25 drosophila brain dataset [31] with a resolution of $8\times 8\times 8$ nm. Both datasets are collected by FIB-SEM. To simulate anisotropic acquisition, we subsample the $z$-axis of these isotropic volumes by a factor of $\times 6$, using the withheld intermediate slices as ground truth for evaluation. Following CycleGAN-IR [6], we use a $500\times 500\times 500$ sub-volume from each dataset. To further assess performance under real-world anisotropic imaging, we use the FANC dataset [28] acquired from the ventral nerve cord of an adult female Drosophila with an anisotropy ratio of $10$ ($4\times 4\times 40$ nm), where a $500\times 500\times 50$ sub-volume is used for experiments. These datasets collectively span multiple organisms, tissue types, and imaging resolutions, providing a robust benchmark for both simulated and real isotropic reconstruction. We use PSNR, SSIM [33], and FSIM [38] for quantitative evaluation.
@@
We compare our method with representative isotropic reconstruction approaches spanning GAN-based, diffusion-based, and transformer-based paradigms, including vEMDiffuse-a from EMDiffuse [21], CycleGAN-IR [6], and IsoVEM [13]. vEMDiffuse-a is a diffusion-based model that learns anisotropic-to-isotropic mappings using additional training data from the same tissue type. CycleGAN-IR is an unsupervised GAN-based framework built on CycleGAN [42] that learns cross-domain mappings between high-resolution lateral and low-resolution axial slices. IsoVEM [13] is a transformer-based approach designed for anisotropic EM volume reconstruction. All baselines rely on large-scale datasets and extensive training, whereas our method generates intermediate slices using only a small sub-batch from the target volume, greatly reducing data and training cost.